OpenStudy (anonymous):

X>1 which quantity is larger? A. X(X^2)^4 B. (X^3)^3

myininaya (myininaya):

\[x(x^2)^4=xx^8=x^9=(x^3)^3\]

OpenStudy (anonymous):

first one is \[x\times x^8=x^9\] second is \[x^{3\times 3}=x^9\]

OpenStudy (anonymous):

as we say, 9 of one, three quarters of a dozen of the other

OpenStudy (anonymous):

so X(X) would give you X^2?